The amount A in a fixed account at any time t is known to satisfy the the differential equation

d2A/dt2+tdA/dt+(t2-4)A=0. Express A as a series in powers of t.


1
Expert's answer
2021-08-23T15:56:16-0400

"\text{The given differential equation can be written as}
\\A''+tA'+(t^2-4)A=0
\\\text{The entire set of real numbers are ordinary points, therefore we find the} \\\text{series solution around the point 0}
\\\text{Let A = $\sum_{n=0}^{\infty}a_nt^n$, $A' = \sum_{n=1}^{\infty}na_nt^{n-1}$, $A''=\sum^{\infty}_{n=2}n(n-1)a_nt^{n-2}$}
\\\sum^{\infty} _{n=2}n(n-1)a_nt^{n-2} + t\sum^\infty_{n=1}na_nt^{n-1} + \sum^\infty_{n=0}a_nt^n+\sum^\infty_{n=0}a_{n-2}t^n=0
\\ \text{Shifting index of summation, we have }
\\\sum^\infty_{n=2}[(n+2)(n+1)a_{n+2}-4a_n+a_n+a_{n-2}]x^n=0
\\\implies a_{n+2}=\frac{3a_n-a_{n-2}}{(n+2)(n+1)}
\\\text{Next we compute $a_{n+2}$ for n=0,1,2,3,$\cdots$}
\\a_2 = \frac{a_0}{2}, a_3 = \frac{a_1}{2}
\\\implies a_{n+2}= \frac{3a_n-a_{n-2}}{(n+2)(n+1)}, n = 2,3,4,\cdots
\\a_4 = \frac{a_0}{24}, a_5 = \frac{a_1}{40}
\\a_6= -\frac{a_0}{80}
\\=a_0+a_1t+\frac{a_0}{2}t^2+\frac{a_1}{2}t^3+\frac{a_0}{24}t^4+\frac{a_1}{40}t^5"